Solve the equation (π‘₯ + 𝑠𝑖𝑛𝑦)𝑑π‘₯ + (𝑦2 + π‘₯π‘π‘œπ‘ π‘¦)𝑑𝑦 = 0


1
Expert's answer
2022-05-24T13:48:28-0400
"\dfrac{\partial Q}{\partial x}=\cos y,\dfrac{\partial P}{\partial y}=\cos y"

We have the following system of differential equations to find the functionΒ  "u(x, y)"


"u=\int(x+\sin y)dx=\dfrac{x^2}{2}+x\sin y+\varphi(y)"

"\dfrac{\partial u}{\partial y}=x\cos y+\varphi'(y)"

"x\cos y+\varphi'(y)=y^2 +x\cos y"

"\varphi'(y)=y^2"

"\varphi(y)=\dfrac{y^3}{3}-C"

So that the general solution of the exact differential equation is given by


"\dfrac{x^2}{2}+x\sin y+\dfrac{y^3}{3}=C"